Charlies Furniture Store has been in business for several years. The firms owners have described the store as a high-price
Answer #1

1. The turnover is 60% of average total assets. So, sales will be $3,000,000*60%= $1,800,000

Return On Investment= Margin* Turnover

= 0.4*0.6

= 0.24 or 24% OR

Return on Investment= Return or profit/ Average total assets

= 720,000/3,000,000

= 24%

2. While being ROI the same and the margin becomes 30% the required turnover would be;

ROI= Margin* Turnover

0.24= 0.3* Turnover

Turnover= 0.24/0.3

= 0.8 or 80%

Required sales amount is; $3,000,000* 80%= $2,400,000

3. The new strategy of the company is to lower the prices, this would reduce the margin from 40% to 30%. The price reduction opted by the company should be accompanied by increase in the sales volume. if not, the strategy would fail. The sales amount required after the new strategy in place while keeping the ROI intact is $ 2,400,000. The current sales is $ 1,800,000. The amount of increase in the sales amount is $ 2,400,000- $ 1,800,000= $ 600,000

4. The new strategy put forward by Charlie to start a new advertising campaign to identify the true and potential customers and directing the marketing efforts to them will be beneficial if made successfully. A successfully placed advertising campaign targeted at the potential customers can definitely bring sales to the company. So, the company can ensure more turnover by not lowering the prices. This can increase the turnover ratio. At the same time the increased advertising and marketing expenses will impact the margin. The more the expenses than the increased sales amount will reduce the margin. There should be a balance between the cost and benefit of undertaking the advertising campaign. The cost of the campaign and the related benefit of increase in sales needs to be worthy. Otherwise the margin will decrease. The decreased margin will affect the Return On Investment. While the investment or total asset remains the same and the margin decreases, the ROI will also decline.

5. Charlie can resort to product differentiation. He could differentiate his products from those offered by the competitors in certain attributes. Depending upon the level of differentiation made, he could charge price differently. He could charge higher price than the competitor if the product offered by him is of superior value than competitor.

He could enter into new markets. He could diversify his marketing and promotional activities to include new market segments. He should attract new customers while retaining the existing customers. He could focus on niche markets.

Other way to remain competitive while protecting his turnover and margin requirements is to lower the costs. He could increase his margin by lowering the cost of production. He should take a complete cost analysis and value chain analysis to determine non-value adding activities and eliminate or reduce them. This can lower the costs or eliminate unwanted costs.This will maintain his ratios in better position.